OpenStudy (anonymous):

5^3t+1=9 i got .12 but im not confident thats the right answer

OpenStudy (anonymous):

\[5^{3t+1}=9\]

OpenStudy (lgbasallote):

seems right

OpenStudy (lgbasallote):

congrats

OpenStudy (lgbasallote):

\[\frac{\log_5 9 - 1}{3} = 0.1217\]